OpenStudy (anonymous):

Help? Solve 2x^2 + 7x - 15 = 0 using the quadratic formula.

OpenStudy (anonymous):

I got [31.25,-32.25].

OpenStudy (jdoe0001):

\(\bf 2x^2 + 7x - 15 = 0 \\ \quad \\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\implies x= \cfrac{ - (7) \pm \sqrt { (7)^2 -4(2)(-15)}}{2(2)}\\ \quad \\ x= \cfrac{ - (7) \pm \sqrt { 49 +120}}{4}\implies x= \cfrac{ - (7) \pm \sqrt { 169}}{4}\implies x= \cfrac{ - 7 \pm 13}{4}\)

OpenStudy (anonymous):

3/2 and - 5

OpenStudy (anonymous):

So [1.5, -20]?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Oops, -5!

OpenStudy (jdoe0001):

yes \(\bf x= \cfrac{ - 7 \pm 13}{4}\\ \quad \\ x = \begin{cases} \cfrac{ - 7 - 13}{4}\implies\cfrac{-20}{4}\implies& \bf -5\\ \quad \\ \bf \cfrac{ - 7 + 13}{4}\implies \cfrac{6}{4}\implies&\bf \cfrac{3}{2} \end{cases}\)

OpenStudy (anonymous):

Thanks!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

I realize now I put a in the place of -(b) instead of 7.

OpenStudy (anonymous):

*2